# Hierarchy Pruning for Unseen Domain Discovery in Predictive Healthcare

### v2

@@ Appendix F Domain Discovery with More Feature Keys
Oracle 16.54 (0.49) 70.12 (0.58) 8.79 (0.50) 68.61 (0.44) 73.11 (0.50) 70.02 (0.47) 67.51 (0.14) 67.28 (0.13)
Base 11.62 (0.60) 55.59 (0.98) 4.21 (0.45) 59.02 (0.73) 50.44 (0.83) 45.03 (0.68) 48.18 (0.21) 45.96 (0.30)
UdonCare 16.03 (0.35) 69.26 (0.40) 6.59 (0.29) 66.45 (0.46) 71.45 (0.33) 67.61 (0.36) 61.81 (0.11) 58.38 (0.23)
@@
Oracle 80.52 (0.13) 67.48 (0.29) 74.62 (0.27) 61.61 (0.23) 26.53 (0.14) 39.57 (0.20) 27.92 (0.12) 40.68 (0.17)
Base 68.39 (0.14) 47.82 (0.34) 66.62 (0.20) 53.47 (0.20) 21.69 (0.15) 31.11 (0.18) 20.22 (0.13) 31.65 (0.20)
UdonCare 78.62 (0.20) 66.12 (0.30) 72.79 (0.35) 59.49 (0.15) 24.55 (0.12) 38.26 (0.21) 27.11 (0.12) 39.63 (0.18)
Appendix G Convergence Analysis via Iterative Learning
@@
Appendix H Upward Information Flow
Considering the hierarchical structure of medical knowledge, a tree-based method (i.e. pruning algorithm) can be naturally adopted for deciding whether we group diseases into higher-level clusters. However, before applying this pruning algorithm to the domain discovery phase, we must ensure that embeddings for all nodes in the hierarchy are available. Since only the leaf-node embeddings $\text{E}(e_{1},\dots,e_{|\mathcal{C}|})$ are obtained from $f_{\phi}(\cdot)$, we propose a unidirectional information flow to initialize and propagate these embeddings upward in a structured yet flexible manner.
First, for each diagnostic code $d_{i}$, we can initialize its embedding $\mathbf{e}_{d_{i}}\in\mathbb{R}^{h}$ in terms of one of two scenarios:
@@ Appendix H Upward Information Flow
We then propagate these embeddings level by level through the hierarchy. The embedding of a parent node $\mathbf{e}^{(h-1)}_{n_{i}}$ is computed from the embeddings of its descendants at level $h$:
$\mathbf{e}^{(h-1)}_{i}\;:=\;\frac{1}{|\mathrm{Desc}(n^{(h-1)}_{i})|}\sum_{n\in\mathrm{Desc}(n_{i})}\mathbf{e}^{(h)}_{n}$ (10)
However, these initial embeddings do not capture the hierarchical distances between node pairs. For example, two codes might have significantly different embeddings despite sharing the same parent node. To address this, and inspired by hierarchical clustering (Johnson, 1967), we rectify these embeddings by considering their relative positions. Following the mechanism of hierarchical clustering, we first calculate cosine similarities of embeddings between pairs $(d_{i},d_{j})$ of leaf nodes or newly formed sub-clusters. The pair with the highest similarity is identified, and their lowest common ancestor $\mathrm{LCA}(d_{i},d_{j})$ in the existing hierarchy is retrieved. The ancestor embedding can be then rectified by incorporating $\mathbf{e}_{d_{i}}$ and $\mathbf{e}_{d_{j}}$:
Any node not identified as an $\mathrm{LCA}$ remains at its average-based embedding from equation 3, and the process repeats until the highest similarity among remaining pairs falls below a predefined threshold. Given the initial embedding table $\text{E}(e_{1},\dots,e_{|\mathcal{C}|})$, this upward flow extends it to $\text{E}(e_{1},\dots,e_{|\mathcal{H}|})$ over the entire hierarchy $\mathcal{H}$. Finally, by incorporating both hierarchical structure and feature similarity through information flow, the resulting node embeddings can capture richer representations upon clinical concept relationships.
Appendix I The Use of Large Language Models (LLMs)

### v3

@@ Appendix H Domain Discovery with More Feature Keys
Oracle 16.95 (0.48) 70.61 (0.58) 8.46 (0.53) 68.92 (0.47) 74.42 (0.43) 69.74 (0.48) 69.37 (0.12) 68.47 (0.11)
Base 11.05 (0.65) 55.01 (0.98) 3.97 (0.47) 59.13 (0.76) 63.45 (0.84) 62.15 (0.73) 58.46 (0.24) 60.56 (0.31)
UdonCare 16.78 (0.35) 70.02 (0.40) 7.21 (0.29) 67.18 (0.46) 72.31 (0.33) 68.32 (0.36) 62.65 (0.11) 59.14 (0.23)
@@
Oracle 80.43 (0.14) 67.41 (0.29) 74.31 (0.25) 61.28 (0.22) 26.97 (0.13) 39.46 (0.27) 28.12 (0.11) 40.53 (0.16)
Base 68.83 (0.19) 57.31 (0.35) 66.94 (0.18) 53.13 (0.18) 21.32 (0.19) 31.08 (0.23) 20.07 (0.12) 31.52 (0.18)
UdonCare 79.34 (0.20) 66.98 (0.30) 73.68 (0.35) 60.22 (0.15) 25.28 (0.12) 38.97 (0.21) 27.86 (0.12) 40.41 (0.18)
Appendix I Convergence Analysis via Iterative Learning
@@
Appendix J Upward Information Flow
To enable hierarchy-aware domain discovery, the model must obtain embeddings for all nodes in the ontology $\mathcal{H}$, not only the leaf concepts $\mathcal{C}$. This appendix details the complete procedure for (i) initializing node features, (ii) propagating them upward through the hierarchy, and (iii) refining ancestor embeddings using hierarchical similarity.
Initializing Leaf-Node Embeddings First, for each diagnostic code $d_{i}$, we can initialize its embedding $\mathbf{e}_{d_{i}}\in\mathbb{R}^{h}$ in terms of one of two scenarios:
@@
Bottom-Up Propagation in the Ontology Since internal nodes have no direct representations, we propagate information from leaf nodes upward. For each parent node $n_{i}^{(h-1)}$ at level h1h-1, its embedding is initialized by averaging the embeddings of its descendants at level $h$:
$\mathbf{e}^{(h-1)}_{i}\;=\;\frac{1}{|\mathrm{Desc}(n^{(h-1)}_{i})|}\sum_{n\in\mathrm{Desc}(n_{i})}\mathbf{e}^{(h)}_{n}$,
which extends the leaf-level table $\text{E}(e_{1},\dots,e_{|\mathcal{C}|})$ to $\text{E}(e_{1},\dots,e_{|\mathcal{H}|})$. However, these initial embeddings do not capture the hierarchical distances between node pairs. For example, two codes might have significantly different embeddings despite sharing the same parent node.
LCA-Based Refinement via Hierarchical Similarity Averaging descendants does not reflect hierarchical distances: two nodes may have similar features even if their parents are far apart in $\mathcal{H}$. To encode such structure, we incorporate a hierarchical-clustering refinement (Johnson, 1967).
(1) Compute pairwise cosine similarity. For any two nodes $(d_{i},d_{j})$ (leaf nodes only), we compute
$\text{cos}(\mathbf{e}_{d_{i}},\mathbf{e}_{d_{j}})=\frac{\mathbf{e}_{d_{i}}^{\top}\mathbf{e}_{d_{j}}}{\|\mathbf{e}_{d_{i}}\|\,\|\mathbf{e}_{d_{j}}\|}.$
Among all possible pairs, we select the one with the highest cosine similarity.
(2) Update the lowest common ancestor (LCA). Let $n_{\mathrm{LCA}}(d_{i},d_{j})$ denote their lowest common ancestor. We update its embedding by averaging itself with the embeddings of the two nodes:
$\mathbf{e}_{\mathrm{LCA}(d_{i},d_{j})}\leftarrow\mathrm{Average}\!\left(\mathbf{e}_{\mathrm{LCA}(d_{i},d_{j})},\mathbf{e}_{d_{i}},\mathbf{e}_{d_{j}}\right).$
(3) Iterative refinement. We repeat this process along the ordered similarity list and stop once the maximum remaining similarity falls below a predefined threshold $\rho=0.3$. Nodes never selected as LCAs retain their initial average-based embeddings from Eq. (3).
Appendix K Scoring Function
The candidate score $S(n)$ involves three indicators:
  1. 1.
    Coverage, $\mathrm{cov}(n)$: measures the ratio of covered leaf nodes $n\in\mathcal{L}$. A higher value indicates greater coverage, favoring the selection of higher-level nodes:
    $\mathrm{cov}(n)=\frac{|\mathcal{M}|}{|\mathcal{L}|},$
    where $\mathcal{M}=\mathrm{Desc}(n)$ denotes the set of descendant leaf nodes of $n$, and $|\mathcal{L}|$ is the total number of leaf nodes in the hierarchy. A higher value indicates that $n$ covers a larger portion of the tree, thus favoring higher-level nodes.
  2. 2.
    Purity, $\mathrm{pur}(n)$: gauges how cohesive the partitioned tree is in the embedding space. This metric evaluates how coherent the subtree rooted at $n$ is in the embedding space. It is computed as the expected similarity between the embedding of node $n$ and the embeddings of its descendant leaf nodes:
    $\mathrm{pur}(n)=\mathbb{E}_{m\in\mathcal{M}}\left[\mathrm{sim}(\mathbf{e}_{n},\mathbf{e}_{m})\right],$
    where $\mathrm{sim}(\cdot)$ denotes the similarity function (e.g., cosine similarity). Higher purity implies that leaf nodes under $n$ are more semantically aligned with the root node.
  3. 3.
    Depth, $\mathrm{dep}(n)\in\{1,\dots,H\}$: indicates the hierarchical level of node $n$. It is defined as:
    $\mathrm{dep}(n)=\frac{h}{H},$
    where $h$ is the depth of node $n$ and $H$ is the maximum depth of the hierarchy. A larger value corresponds to finer-grained concepts, thus favoring deeper nodes.
Appendix L Broader Applicability and Extensions of UdonCare
While UdonCare is evaluated primarily on structured, code-based EHR data, the core design of the framework is modality-agnostic and can be extended beyond the four clinical event prediction tasks explored in this work. We briefly highlight two promising directions:
L.1 Extension to Multi-modal Clinical Data
The ontology-guided domain discovery mechanism is compatible with richer, multi-modal patient records. Modern hospital systems routinely collect diverse modalities, including:
  1. 1.
    Clinical notes (text), encoded with contextualized language models;
  2. 2.
    Radiology and pathology images, structured under well-defined ontologies such as RadLex;
  3. 3.
    Physiological waveforms (ECG, EEG), which contain temporally structured signals;
  4. 4.
    Laboratory trajectories, which follow standardized concept hierarchies (LOINC).
In principle, our model can incorporate these modalities by (1) extending the feature extractor to relevant encoders (e.g., CNNs, transformers for images, waveform encoders), and (2) pruning cross-modal ontologies or multi-level taxonomies whenever available. The mutual-learning mechanism in UdonCare, designed to decouple domain and label information, benefits from such multi-modal richness by preventing modality-specific biases from dominating the latent domain assignments. However, large-scale multi-modal extensions require access to paired datasets, which is unavailable in MIMIC or other standard public EHR databases. Thus, multi-modal generalization remains an important but data-dependent direction for future research.
L.2 Generalization Beyond Healthcare to Ontology-Rich Domains
The ontology-guided design is also not restricted to clinical settings. Many real-world domains exhibit hierarchical or taxonomy-driven feature structures that mirror the ICD hierarchy, such as
  1. 1.
    E-commerce: product category graphs, brand hierarchies, user segments;
  2. 2.
    Recommendation systems: taxonomies of items or content genres;
  3. 3.
    Knowledge-driven applications such as academic topic hierarchies, legal codes, or biological ontologies.
In these domains, domain shifts commonly arise from evolving user groups, temporal drifts in product popularity or region-specific differences, which are analogous to the clinical temporal and spatial shifts studied in this work. The pruning-based latent domain discovery can meaningfully identify stable high-level clusters across such hierarchies, while the invariant-projection mechanism can separate domain-specific variations from task-relevant signals. The adaptability of its ontology-guided components suggests broad potential applicability to many hierarchical, knowledge-grounded machine learning scenarios.
Appendix M More Explanations on the Learned Linear Weights
In this section, we discuss the relatively low cosine similarity in Table 2:
First, a similar trend has been observed in ManyDG (Yang et al., 2023a). Linear probing for domain prediction is inherently more challenging than probing for label prediction. The domain classification task usually contains a substantially larger number of categories (e.g. over 500), whereas the label space is considerably smaller. Thus, the learned linear coefficients for domain prediction tend to be less aligned, resulting in lower cosine similarity in the second row.
Second, if the model successfully decomposes patient embeddings into label-related and domain-related components, the observed similarity pattern becomes intuitive. The first row corresponds to label prediction, and thus naturally shows higher similarity because label information constitutes a larger and stronger portion of the embedding space. In contrast, after label (domain-invariant) information is accounted for, the remaining domain signal becomes weaker, making it harder for a linear classifier to capture consistently. This yields lower similarity in the second row.
We emphasize that this explanation does not imply a precise quantitative ratio between label and domain information. Rather, the key observation is that when the invariant (label-relevant) component dominates the embedding space, the remaining domain-specific signal becomes harder to recover linearly, and this phenomenon is reflected consistently in our empirical results.
Appendix N The Use of Large Language Models (LLMs)